# ATTS: Asynchronous Test-Time Scaling via Conformal Prediction

### v2

@@ Abstract
Large language models (LLMs) benefit from test-time scaling but are often hampered by high inference latency. Speculative decoding is a natural way to accelerate the scaling process; however, scaling along both the parallel and sequential dimensions poses significant challenges, including substantial memory-bound execution and synchronization overhead. We introduce ATTS (Asynchronous Test-Time Scaling), a statistically guaranteed adaptive scaling framework that follows the hypothesis testing process to address these challenges. By revisiting arithmetic intensity, ATTS identifies synchronization as the primary bottleneck. It enables asynchronous inference through online calibration and proposes an ordinal classification algorithm that supports a three-stage rejection sampling pipeline, scaling along both the sequential and parallel axes. Across experiments on the MATH, AMC23, AIME24, and AIME25 datasets and across multiple draft–target model families, we show that ATTS delivers up to 56.7x speedup in test-time scaling and a 4.14x throughput improvement, while maintaining accurate control of the rejection rate, reducing latency and memory overhead, and incurring no accuracy loss. By scaling both in parallel and sequential dimensions, we enable the 1.5B/70B draft/target model combination to achieve the performance of the state-of-the-art reasoning model o3-mini (high) on the AIME dataset. We have released the code at https://github.com/menik1126/asynchronous-test-time-scaling.
@@ 1 Introduction
With the rapid advances in large language models (LLMs), attention is increasingly turning to reasoning models (Guo et al., 2025; Muennighoff et al., 2025; McCoy et al., 2024; Shao et al., 2024)—systems that transcend next-token prediction in order to emulate human-like reasoning behaviors. These models excel at leveraging complex reasoning chains, especially in test-time scaling settings (Snell et al., 2024; Li et al., 2025; Muennighoff et al., 2025; Zeng et al., 2025).
@@
Speculative decoding (Li et al., 2024a; Leviathan et al., 2023; Kim et al., 2023; Pan et al., 2025b; Yang et al., 2025) represents a promising approach for accelerating decoding. In this framework, a lightweight draft model generates tokens, which are subsequently validated and refined by a target model. This dual-phase approach not only speeds up inference by offloading most of the generation process to the draft model, but also ensures that the final outputs retain high fidelity, thereby achieving a favorable balance between efficiency and accuracy.
@@ 3 Challenge and Design
To identify performance bottlenecks in the classic setup, we introduce arithmetic intensity (Spector & Re, 2023), which measures the utilization of arithmetic units. It is defined as:
@@ 3.2 Q2: How is the prediction set constructed?
Conformal prediction typically relies on a held-out calibration set to determine the threshold $\tau$. However, in the test-time scaling setup, held-out examples are generally unavailable. To address this limitation, we propose an online calibration strategy. Specifically, $m$ outputs are pre-sampled for each input in the test set, yielding $(\hat{Y}_{i}^{1},\dots,\hat{Y}_{i}^{m})$. We then relax the strict requirement that conformity scores sum to one, and instead directly define:
@@
Let $B$ denote the predefined budget (i.e., the number of candidates to reject). Given a test-time input $X_{i}$, we sample $m$ candidate CoTs $\big(\hat{Y}_{i}^{1},\dots,\hat{Y}_{i}^{m}\big)$ in each turn and then compute their corresponding $p$-values $p_{i}^{1},\dots,p_{i}^{m}$}.
@@
$\mathcal{C}_{\alpha}(Y_{i})=\left\{\hat{Y}_{i}^{k}:k\in\{1,\ldots,m\},p_{i}^{k}>\alpha\right\}.$ (15)
@@ 3.3 How to perform rejection sampling via conformal prediction?
For each candidate sampling $\tilde{y}_{n:n+K_{d}-1}^{j}$, we score it under the target model $q_{t}$ by computing the logits $q_{t}(\tilde{y}_{n:n+K_{d}-1}^{j}\mid x_{1:n-1})$ and converting to a conformity score. Using the calibration set, we compute a $p$-value for each candidate and reject those inside $\mathcal{C}_{\alpha}$; otherwise accept.
@@
Although classical rejection sampling discards the rejected samples from the draft model and resamples entirely from the target model, to save the token budget we proceed as follows. In each turn, the per-turn target-side token budget is $K_{t}$: for each candidate in $\mathcal{C}_{\alpha}$, we let the target model $q_{t}$ continue generation using that candidate from $q_{d}$ as a prefix for up to $K_{t}$ tokens, stopping earlier if an end token is encountered. In Appendix A.4, we provide a comparison between continuing sampling and resampling with $q_{t}$ for large-scale scaling performance.
@@ 4 Experiment
Table 2: The comparison when the draft and target models are from the same family. $\bm{T^{m}_{\text{Think}}}$, $\bm{T^{m}_{\text{Reason}}}$, $\bm{T^{c}_{\text{Think}}}$, and $\bm{T^{c}_{\text{Reason}}}$ denote the token-consumption ratios (×) under the marginal coverage and conditional coverage settings, relative to the SpecThink and SpecReason baselines, respectively. Token consumption is measured in tokens, where smaller values indicate higher efficiency.
Dataset
Draft Model / Target Model
Mar Cov.
Con Cov.
SpecThink
SpecReason
Token Consumption (×)
$\bm{T^{m}_{\text{Think}}}$
$\bm{T^{m}_{\text{Reason}}}$
$\bm{T^{c}_{\text{Think}}}$
$\bm{T^{c}_{\text{Reason}}}$
@@
Figure 6: The budget prediction accuracy shows the accuracy of budget prediction for sampling lengths of 500 and 700 with a budget of 500 tokens during the online calibration phase under 64-sample and 16-sample settings. The experiment is conducted with a rejection rate $\alpha=0.25$.
@@ 4.4 Analysis of token budget and latency
Figure 7 presents the token consumption under the 16-sample setting. This includes sampling performed solely by the target model or the draft model as baselines, as well as asynchronous sampling under both condition coverage and marginal coverage settings. Compared with the two baselines, our method can significantly reduce token consumption, especially under the condition coverage setting, as it enables budget prediction at the instance level.
@@ A.2 Related Work
Speculative decoding(Li et al., 2024a; Leviathan et al., 2023; Kim et al., 2023; Chen et al., 2023) is an emerging technique for accelerating LLM inference, which is traditionally limited by slow, sequential autoregressive sampling and memory bandwidth constraints. There are three main strategies for sampling draft tokens: token-level sampling(Leviathan et al., 2023; Kim et al., 2023; Chen et al., 2023), where the large model directly verifies the token outputs of the draft model; feature-level sampling (Cai et al., 2024; Li et al., 2024a; b), which verifies generation paths using intermediate representations; and step-level sampling (Pan et al., 2025b; Yang et al., 2025), which operates at a coarser granularity by validating multiple tokens or computation steps together to improve throughput.
@@
A.3 Experimental Setup
@@ A.5 Ablation Study
A.6 Controlling the rejection rate in multi-turn interactions
@@ Discussion.
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in\mathcal{C}(X_{i}))\geq 1-\alpha.$ (35)
@@
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in\mathcal{C}(X_{i}))\geq 1-\alpha+\frac{1}{(n-1)m+1}.$ (36)
@@
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in\mathcal{C}(X_{i}))$ $=\mathbb{P}(\text{accept }H_{Y_{\xi}})$ (37)
@@
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in\mathcal{C}(X_{i}))$ $=\mathbb{P}(\text{accept }H_{Y_{\xi}})$ (41)
@@
Suppose that $\{(X_{i},Y_{i})\}_{i=1}^{n}$ are exchangeable random variables, and let $\xi\sim\text{Uniform}\{1,2,\ldots,n\}$ represent a randomly selected test instance, then the prediction set $\mathcal{C}(X_{\xi})=\left\{\hat{Y}_{\xi}^{k}\mid p_{\xi}^{k}>\alpha\right\}$ determined by A1 both satisfy
$\mathbb{P}(Y_{\xi}\in\mathcal{C}(X_{\xi}))\geq 1-\alpha.$ (45)
@@
Note that the prediction set is given by $\mathcal{C}(X_{\xi})=A_{1}\cap A_{2}$. Thus, by Proposition 1,
$\mathbb{P}(Y_{\xi}\in\mathcal{C}(X_{\xi}))\geq\mathbb{P}(p_{\xi}^{Y_{\xi}}>\alpha)\geq 1-\alpha.$ (46)
@@
$\mathbb{P}(Y_{\xi}\in\mathcal{C}(X_{\xi}))=\mathbb{P}(p_{\xi}^{Y_{\xi}}>\alpha)\geq 1-\alpha.$ (48)
@@
$\mathbb{P}(Y_{\xi}\in\mathcal{C}(X_{\xi}))=1-\mathbb{P}(p_{\xi}^{Y_{\xi}}\leq\alpha)\leq 1-\alpha+\frac{1}{(n-1)m+1}.$ (49)
@@
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in\mathcal{C}(X_{i})\mid Y_{\xi}=y)\geq 1-\alpha.$ (50)
@@
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in\mathcal{C}(X_{i})\mid Y_{\xi}=y,\mathcal{I}_{y})\geq 1-\alpha+\frac{1}{(n-1)|\mathcal{I}_{y}|+1}.$ (51)
@@
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in\mathcal{C}(X_{i})\mid Y_{\xi}=y)$ $=1-\mathbb{P}(\text{reject }H_{y}\mid Y_{\xi}=y)$ (52)
@@
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in\mathcal{C}(X_{i})\mid\mathcal{I}_{y},Y_{\xi}=y)$ $=1-\mathbb{P}(\text{reject }H_{y}\mid\mathcal{I}_{y},Y_{\xi}=y)$ (55)
@@
Under the same exchangeability assumption as in Theorem 1, the prediction set $\mathcal{C}(X_{\xi}\mid y)$ based on conditional conformal $p$-values $p_{\xi}^{k}$ satisfies
$\mathbb{P}(Y_{\xi}\in\mathcal{C}(X_{\xi}\mid y)\mid Y_{\xi}=y)\geq 1-\alpha$ (59)
for any $y\in\mathcal{Y}$. Specifically, for the prediction set $\mathcal{C}(X_{\xi}\mid y)$ based on $p_{\xi}^{k}$, if the conformity scores $\{s_{i}^{j}\}_{j\in\mathcal{I}_{y},i=1,\ldots,n}$ are distinct surely, we have
$\mathbb{P}(Y_{\xi}\in\mathcal{C}(X_{\xi}\mid y)\mid Y_{\xi}=y,\mathcal{I}_{y})\leq 1-\alpha+\frac{1}{(n-1)|\mathcal{I}_{y}|+1}$ (60)
@@
By using Proposition 4 and similar arguments as in the proof of Theorem 1, the prediction sets $\mathcal{C}(X_{\xi}\mid y)$ based on the conditional conformal $p$-values $p_{\xi}^{k}$ all satisfy:
$\mathbb{P}(Y_{\xi}\in\mathcal{C}(X_{\xi}\mid y)\mid\mathcal{I}_{y},Y_{\xi}=y)\geq 1-\alpha$ (61)
@@
$\mathbb{P}(Y_{\xi}\in\mathcal{C}(X_{\xi}\mid y)\mid\mathcal{I}_{y},Y_{\xi}=y)$ $=\mathbb{P}(p_{\xi}^{y}>\alpha\mid\mathcal{I}_{y},Y_{\xi}=y)$ (62)

### v3

@@ Abstract
Large language models (LLMs) benefit from test-time scaling but are often hampered by high inference latency. Speculative decoding is a natural way to accelerate the scaling process; however, scaling along both the parallel and sequential dimensions poses significant challenges, including substantial memory-bound execution and synchronization overhead. We introduce ATTS (Asynchronous Test-Time Scaling), a statistically guaranteed adaptive scaling framework that follows the hypothesis testing process to address these challenges. By revisiting arithmetic intensity, ATTS identifies synchronization as the primary bottleneck. It enables asynchronous inference through online calibration and proposes an ordinal classification algorithm that supports a three-stage rejection sampling pipeline, scaling along both the sequential and parallel axes. Across experiments on the MATH, AMC23, AIME24, and AIME25 datasets and across multiple draft–target model families, we show that ATTS delivers up to 56.7x speedup in test-time scaling and a 4.14x throughput improvement, while maintaining accurate control of the rejection rate, reducing latency and memory overhead, and incurring no accuracy loss. By scaling both in parallel and sequential dimensions, we enable the 1.5B/70B draft/target model combination to achieve the performance of the state-of-the-art reasoning model o3-mini (high) on the AIME dataset.
@@ 1 Introduction
With the rapid advances in large language models (LLMs), attention is increasingly turning to reasoning models (Guo et al., 2025; Muennighoff et al., 2025; McCoy et al., 2024; Shao et al., 2024)—systems that transcend next-token prediction in order to emulate human-like reasoning behaviors. These models excel at leveraging complex reasoning chains, especially in test-time scaling settings (Snell et al., 2024; Li et al., 2025; Muennighoff et al., 2025; Zeng et al., 2025), and have shown strong potential in mathematical reasoning (Xiong et al., 2022; 2023b; 2023a).
@@
Speculative decoding (Li et al., 2024b; Leviathan et al., 2023; Kim et al., 2023; Pan et al., 2025b; Yang et al., 2025) represents a promising approach for accelerating decoding. In this framework, a lightweight draft model generates tokens, which are subsequently validated and refined by a target model. This dual-phase approach not only speeds up inference by offloading most of the generation process to the draft model, but also ensures that the final outputs retain high fidelity, thereby achieving a favorable balance between efficiency and accuracy.
@@ 3 Challenge and Design
To identify performance bottlenecks in the classic setup, we introduce arithmetic intensity (Spector and Re, 2023), which measures the utilization of arithmetic units. It is defined as:
@@ 3.2 Q2: How is the prediction set constructed?
Conformal prediction typically relies on a held-out calibration set to determine the threshold $\tau$. However, in the test-time scaling setup, held-out examples are generally unavailable. To address this limitation, we propose an online calibration strategy. Specifically, $m$ outputs are pre-sampled for each input in the test set, yielding $(\hat{Y}_{i}^{1},\dots,\hat{Y}_{i}^{m})$. Previous efforts (Ding et al., 2023; Romano et al., 2020) impose a strict sum-to-one constraint on the conformal scores under the classification setting (where events are mutually exclusive). In contrast, we compute conformal p-values (Bates et al., 2023; Jin and Candès, 2023; Wang et al., 2024) under an ordinal classification setup, where the events are not mutually exclusive and the ordinal relationships are preserved. In this setup, we relax the strict requirement that conformity scores sum to one, and instead directly define:
@@
Let $B$ denote the predefined budget (i.e., the number of candidates to reject). Given a test-time input $X_{\xi}$, we sample $m$ candidate CoTs $\big(\hat{Y}_{\xi}^{1},\dots,\hat{Y}_{\xi}^{m}\big)$ in each turn and then compute their corresponding $p$-values $p_{\xi}^{1},\dots,p_{\xi}^{m}$}.
@@
$C_{\alpha}(Y_{\xi})=\left\{\hat{Y}_{\xi}^{k}:k\in\{1,\ldots,m\},p_{\xi}^{k}>\alpha\right\}.$ (15)
@@ 3.3 How to perform rejection sampling via conformal prediction?
For each candidate sampling $\tilde{y}_{N:N+K_{d}-1}^{j}$, we score it under the target model $q_{t}$ by computing the logits $q_{t}(\tilde{y}_{N:N+K_{d}-1}^{j}\mid x_{1:N-1})$ and converting to a conformity score. Using the calibration set, we compute a $p$-value for each candidate and reject those inside $C_{\alpha}$; otherwise accept.
@@
Although classical rejection sampling discards the rejected samples from the draft model and resamples entirely from the target model, to save the token budget we proceed as follows. In each turn, the per-turn target-side token budget is $K_{t}$: for each candidate in $C_{\alpha}$, we let the target model $q_{t}$ continue generation using that candidate from $q_{d}$ as a prefix for up to $K_{t}$ tokens, stopping earlier if an end token is encountered. In Appendix A.5, we provide a comparison between continuing sampling and resampling with $q_{t}$ for large-scale scaling performance.
@@ 4 Experiment
Table 2: Comparison within the same model family. $\bm{T^{m}_{\text{Th}}}$$\bm{T^{m}_{\text{Re}}}$ and $\bm{T^{c}_{\text{Th}}}$$\bm{T^{c}_{\text{Re}}}$ denote token-consumption ratios ($\times$) under marginal/conditional coverage relative to SpecThink (Th) and SpecReason (Re). Values in red indicate lossless acceleration (accuracy $\geq$ TM baseline). Gray rows: Skywork-OR1 draft model.
Accuracy Token Consumption ($\times$)
Dataset Draft / Target Model Mar. Cond. SpecTh. SpecRe. $T^{m}_{\text{Th}}$ $^{m}_{\text{Re}} TT$$^{c}_{\text{Th}}$ $T^{c}_{\text{Re}}$
@@
051015202530Value01234567891011121314Problem64 samples predict 50064 samples predict 70016 samples predict 50016 samples predict 700
@@
Figure 6: Budget prediction accuracy with a rejection rate $\alpha=0.25$. (a) Per-batch conditional coverage: each bar represents the prediction error per problem under different settings ($K$ samples predict $L$ denotes $K$ parallel samples with calibration budget 500 and sampling budget $L$). (b) Dataset-level marginal coverage: Take Over denotes the fraction of samples handled by the target model, and Remaining denotes those kept from the draft model.
@@ 4.4 Analysis of token budget and latency
Figure 8(a) presents the token consumption under the 16-sample setting. This includes sampling performed solely by the target model or the draft model as baselines, as well as asynchronous sampling under both condition coverage and marginal coverage settings. Compared with the two baselines, our method can significantly reduce token consumption, especially under the condition coverage setting, as it enables budget prediction at the instance level.
4.5 Multi-turn Evaluation Results
In this section, we evaluate the results on the AIME25 dataset under settings with more turns. Unlike the previous setting, we reduce the token budget per turn but increase the number of iterations, allowing the target model to generate more samples. As shown in Figure 8(b), the accuracy increases with the number of turns. Both marginal coverage and conditional coverage eventually converge to the same value. The results show that under a fixed sampling size budget, increasing the number of turns still does not break the performance upper bound.
020004000600080001000012000Token BudgetAIME24AIME25AMC23MATH100DatasetTMTM Cond.TM Marg.DMDM Cond.DM Marg.
(a) Token consumption under 16-sample setting. TM/DM = Target/Draft Model; Cond./Marg. = Conditional/Marginal Coverage.
0102030405060Accuracy (%)2481216202428TurnDeepSeek Mar.DeepSeek Cond.Qwen Mar.Qwen Cond.
(b) Multi-turn evaluation on AIME25. Solid/dashed lines denote Mar./Cond. coverage. DeepSeek: 1.5B/S1.1-32B; Qwen: 7B/S1.1-32B.
Figure 8: (a) Token consumption under the 16-sample setting across different datasets. (b) Multi-turn evaluation results on AIME25 with increasing turns.
@@ 4.5 Multi-turn Evaluation Results
Acknowledgments
This work was supported in part by the Theme-based Research Scheme (TRS) project T45-701/22-R of the Research Grants Council of Hong Kong, and in part by the AVNET-HKU Emerging Microelectronics and Ubiquitous Systems (EMUS) Lab.
@@ A.2 Related Work
Speculative decoding(Li et al., 2024b; Leviathan et al., 2023; Kim et al., 2023; Chen et al., 2023) is an emerging technique for accelerating LLM inference, which is traditionally limited by slow, sequential autoregressive sampling and memory bandwidth constraints. There are three main strategies for sampling draft tokens: token-level sampling(Leviathan et al., 2023; Kim et al., 2023; Chen et al., 2023), where the large model directly verifies the token outputs of the draft model; feature-level sampling (Cai et al., 2024; Li et al., 2024b; a), which verifies generation paths using intermediate representations; and step-level sampling (Pan et al., 2025b; Yang et al., 2025), which operates at a coarser granularity by validating multiple tokens or computation steps together to improve throughput.
@@
A.3 OlympiadBench
We evaluate the results on the more challenging OlympiadBench (He et al., 2024) dataset as evidence of the robustness of our method. Under two of our settings (results shown in bold), we were even able to surpass the performance of the original target model’s sampling. Under conditional coverage, we achieved a more efficient allocation of computational resources compared to marginal coverage, resulting in improved test performance. Table 3 shows the results on the OlympiadBench dataset.
Table 3: Results on the OlympiadBench benchmark. Each setting uses 16 samples and 15 turns, with $\alpha=0.4$, 500-token budgets, and temperature 0.8.
Draft / Target Model Draft Model Target Model Marginal Cov. Conditional Cov.
DeepSeek-R1-Distill-Llama-8B / S1.1-32B 28 48 44 40
DeepSeek-R1-Distill-Qwen-1.5B / S1.1-32B 26 48 38 38
Llama-3.1-8B-Instruct / S1.1-32B 26 48 44 50
Qwen2.5-7B-Instruct / S1.1-32B 32 48 48 48
DeepSeek-R1-Distill-Llama-8B / QwQ-32B 28 46 38 40
DeepSeek-R1-Distill-Qwen-1.5B / QwQ-32B 26 46 38 40
Llama-3.1-8B-Instruct / QwQ-32B 26 46 48 48
Qwen2.5-7B-Instruct / QwQ-32B 32 46 50 46
A.4 Experimental Setup
@@ A.6 Ablation Study
A.7 Controlling the rejection rate in multi-turn interactions
@@ Discussion.
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in C(X_{i}))\geq 1-\alpha.$ (35)
@@
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in C(X_{i}))\geq 1-\alpha+\frac{1}{(n-1)m+1}.$ (36)
@@
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in C(X_{i}))$ $=\mathbb{P}(\text{accept }H_{Y_{\xi}})$ (37)
@@
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in C(X_{i}))$ $=\mathbb{P}(\text{accept }H_{Y_{\xi}})$ (41)
@@
Suppose that $\{(X_{i},Y_{i})\}_{i=1}^{n}$ are exchangeable random variables, and let $\xi\sim\text{Uniform}\{1,2,\ldots,n\}$ represent a randomly selected test instance, then the prediction set $C(X_{\xi})=\left\{\hat{Y}_{\xi}^{k}\mid p_{\xi}^{k}>\alpha\right\}$ determined by A1 both satisfy
$\mathbb{P}(Y_{\xi}\in C(X_{\xi}))\geq 1-\alpha.$ (45)
@@
Note that the prediction set is given by $C(X_{\xi})=A_{1}\cap A_{2}$. Thus, by Proposition 1,
$\mathbb{P}(Y_{\xi}\in C(X_{\xi}))\geq\mathbb{P}(p_{\xi}^{Y_{\xi}}>\alpha)\geq 1-\alpha.$ (46)
@@
$\mathbb{P}(Y_{\xi}\in C(X_{\xi}))=\mathbb{P}(p_{\xi}^{Y_{\xi}}>\alpha)\geq 1-\alpha.$ (48)
@@
$\mathbb{P}(Y_{\xi}\in C(X_{\xi}))=1-\mathbb{P}(p_{\xi}^{Y_{\xi}}\leq\alpha)\leq 1-\alpha+\frac{1}{(n-1)m+1}.$ (49)
@@
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in C(X_{i})\mid Y_{\xi}=y)\geq 1-\alpha.$ (50)
@@
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in C(X_{i})\mid Y_{\xi}=y,\mathcal{I}_{y})\geq 1-\alpha+\frac{1}{(n-1)|\mathcal{I}_{y}|+1}.$ (51)
@@
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in C(X_{i})\mid Y_{\xi}=y)$ $=1-\mathbb{P}(\text{reject }H_{y}\mid Y_{\xi}=y)$ (52)
@@
$\mathbb{P}(\forall i\in\{1,\ldots,n\}:Y_{i}\in C(X_{i})\mid\mathcal{I}_{y},Y_{\xi}=y)$ $=1-\mathbb{P}(\text{reject }H_{y}\mid\mathcal{I}_{y},Y_{\xi}=y)$ (55)
@@
Under the same exchangeability assumption as in Theorem 1, the prediction set $C(X_{\xi}\mid y)$ based on conditional conformal $p$-values $p_{\xi}^{k}$ satisfies
$\mathbb{P}(Y_{\xi}\in C(X_{\xi}\mid y)\mid Y_{\xi}=y)\geq 1-\alpha$ (59)
for any $y\in\mathcal{Y}$. Specifically, for the prediction set $C(X_{\xi}\mid y)$ based on $p_{\xi}^{k}$, if the conformity scores $\{s_{i}^{j}\}_{j\in\mathcal{I}_{y},i=1,\ldots,n}$ are distinct surely, we have
$\mathbb{P}(Y_{\xi}\in C(X_{\xi}\mid y)\mid Y_{\xi}=y,\mathcal{I}_{y})\leq 1-\alpha+\frac{1}{(n-1)|\mathcal{I}_{y}|+1}$ (60)
@@
By using Proposition 4 and similar arguments as in the proof of Theorem 1, the prediction sets $C(X_{\xi}\mid y)$ based on the conditional conformal $p$-values $p_{\xi}^{k}$ all satisfy:
$\mathbb{P}(Y_{\xi}\in C(X_{\xi}\mid y)\mid\mathcal{I}_{y},Y_{\xi}=y)\geq 1-\alpha$ (61)
@@
$\mathbb{P}(Y_{\xi}\in C(X_{\xi}\mid y)\mid\mathcal{I}_{y},Y_{\xi}=y)$ $=\mathbb{P}(p_{\xi}^{y}>\alpha\mid\mathcal{I}_{y},Y_{\xi}=y)$ (62)